If 165.7 MJ/mol photon strike atoms of an unknown element and eject electrons with kinetic energies of 25.4 MJ/mol, what is the workfunction of the element? 

What is the frequency of light attributed to the n=3 line in the Balmer transition (n=2) of hydrogen emission


1
Expert's answer
2020-12-06T17:21:10-0500

Solution

Using Einstein equation

"E=\phi+KE"

So work function

"\phi=E-KE=165.7-25.4\\=140.3MJ/mol"


Wavelength is given by

"\frac{1}{\lambda}=R(\frac{1}{n^2_f}-\frac{1}{n^2_i})"

Here

"n_i=3\\n_f=2" putting all value in above equation

Wavelength is found

"\lambda=656.88nm"

So frequency is found

"\nu=\frac{c}{\lambda}=\frac{3\times10^8}{656.88nm}\\=4567\times 10^{11}Hz"